OpenStudy (anonymous):

Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation. 27.776 g-Measured mass of metal 26 mL- Distilled water measurement 25 C- Distilled water temp 100 C- Temp of metal 31.7 C- temp of mixture

OpenStudy (phi):

can you figure out q_water ? using m= 26 g , c= 4.18, and ΔT=(31.7-25)

OpenStudy (anonymous):

i'm sorry what do you mean by that?

OpenStudy (anonymous):

how would I find that out?

OpenStudy (phi):

your formula qwater = m × c × ΔT. says: multiply 3 things together: the mass, c (which is 4.18 because they told you) and the change in temperature. The water started at 25 C and went up to 31.7C

OpenStudy (anonymous):

ok hold on right quick

OpenStudy (anonymous):

it would be 728.156, that would be the qwater correct?

OpenStudy (phi):

yes. now for the metal they do not tell you c (that is what you have to find) but you know the rest of the info

OpenStudy (anonymous):

@phi : hold on, I thought you just told me what c was?

OpenStudy (anonymous):

@phi: which was 4.18?

OpenStudy (phi):

c is the specific heat capacity of a material. It is different for different things. they want you to find c for the metal

OpenStudy (anonymous):

@phi: well they also give us a unknown metal that we had to calculate. This is the unknown metal table: It goes in the same order. 25.605g 24 mL 25 C 100 C 0 C

OpenStudy (anonymous):

@phi : they then tell us to do the same. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qunknown metal Using the formula qunknown metal = m × c × ΔT, calculate the specific heat of the metal.

OpenStudy (phi):

ok, but first let's finish your first problem

OpenStudy (phi):

Your first problems says (for the metal) 27.776 g-Measured mass of metal 100 C- Temp of metal 31.7 C- temp of mixture It looks like the metal was hot, and they put it into water. the water heated up (the water absorbed the heat 728.156 Joules according to your calculation) the metal gave off that amount of heat and cooled down

OpenStudy (phi):

the idea is that qmetal = m × c × ΔT but q_metal (the amount of heat the metal lost), equals the amount of heat that went into the water. so we can say q_metal= 728.156 Joules

OpenStudy (phi):

now fill in as much as you know in the formula qmetal = m × c × ΔT (you know everything except "c" , which is the specific heat of this metal)

OpenStudy (anonymous):

so then to fill in the formula it would be 728.156 Joules= 26 g (x) c (x) 6.7

OpenStudy (anonymous):

@phi: Am I correct?

OpenStudy (phi):

that is for the water (and c for water is 4.18 ) we want to do the same thing for the metal

OpenStudy (anonymous):

ok now can you explain to me how you got that 4.18 so that I can try to see how I would get c for the metal?

OpenStudy (phi):

They told us c for water is 4.18 ***the specific heat capacity of water is 4.18 J / (g × °C) *** we used the formula q_water = m × c × ΔT to find q_water= 728.156 Joules The idea is that heat is energy. It took 728.156 J of energy to heat up the water. where did that energy come from ? It came from the metal. The metal "lost" 728.156 J of energy If we use the formula q_metal = m × c × ΔT replace q_metal with 728.156 J also, fill in the mass of the metal, and how much the temperature of the metal changed.

OpenStudy (anonymous):

Ok so now it is for metal: 728.156 J= 25.605 x c x (0-25) Correct?

OpenStudy (phi):

I think we should stick to the first problem you posted. Use the data from your experiment for the metal in your calculation. 27.776 g-Measured mass of metal 26 mL- Distilled water measurement 25 C- Distilled water temp 100 C- Temp of metal 31.7 C- temp of mixture

OpenStudy (anonymous):

ok we are on the first problem, we are back to it, a clean slate. We have just figured out the Joules, now what would be the next step?

OpenStudy (phi):

q_metal = m × c × ΔT replace q_metal with 728.156 J also, fill in the mass of the metal, and how much the temperature of the metal changed.

OpenStudy (anonymous):

ok wait so now we are on the 2nd question?

OpenStudy (phi):

This is the question you posted. We have done the water part and got 728.156 J now we are doing the metal

OpenStudy (anonymous):

oh now I get it, we are doing the =qmetal correct?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok so now that I get it lol how would be find c?

OpenStudy (phi):

The next step is q_metal = m × c × ΔT replace q_metal with 728.156 J also, fill in the mass of the metal, and how much the temperature of the metal changed.

OpenStudy (anonymous):

So it is now 728.156 J= 27.776 x 75?

OpenStudy (phi):

how did you get 75 ? (and be sure to leave the c in the equation)

OpenStudy (anonymous):

100 degrees C is the temp of the metal and 25 degrees C is the distilled water temp

OpenStudy (phi):

I think the 25 C is the water before you put the metal into the water. after a little while, the water heated up to 31.7 (and the metal cooled down to 31.7) Does that make sense ?

OpenStudy (anonymous):

no it doesn't.... sorry :(

OpenStudy (phi):

if you have some water and put a chunk of hot metal into the water, what happens to the temperature of the water?

OpenStudy (anonymous):

the temp would get hot

OpenStudy (phi):

and the metal would cool down until the water and the metal reached the same temperature. That is what is going on here. when they tell us 25 C- Distilled water temp 100 C- Temp of metal 31.7 C- temp of mixture the 25 C is the water and the 100 C is the hot metal 31.7 C is the temperature of the metal in the water (the mixture) after the water heated up and the metal cooled down until both are at the same temperature. (notice that when you mix two things together they reach the *same* temperature... it would be weird if you put hot water in your cup and the cup got colder and the water got hotter... it doesn't do that... they water and the cup come to the same temperature)

OpenStudy (anonymous):

oh ok I get that analogy. thanks so now what will happen next.

OpenStudy (phi):

The next step is q_metal = m × c × ΔT replace q_metal with 728.156 J also, fill in the mass of the metal, and how much the temperature of the metal changed.

OpenStudy (anonymous):

ok so now it would be the mixture correct?

OpenStudy (phi):

for the temperature (both the water and the metal reach the same temperature)

OpenStudy (anonymous):

ok so it would be 728.156= 27.776 x 31.7?

OpenStudy (phi):

two things: (1) be sure to KEEP the "c" (2) ΔT is the change in the temperature of the metal. How many degrees did it change ?

OpenStudy (anonymous):

ok so it is 728.156= 27.776 x c x ? I don't get it when you say how many degrees did it change.

OpenStudy (phi):

read this carefully 25 C- Distilled water temp 100 C- Temp of metal 31.7 C- temp of mixture the 25 C is the water and the 100 C is the hot metal 31.7 C is the temperature of the metal in the water (the mixture) after the water heated up and the metal cooled down

OpenStudy (anonymous):

is the answer 31.7?

OpenStudy (anonymous):

since the temp becomes the same?

OpenStudy (phi):

31.7 C is the temperature of the metal after it cooled down. but it cooled down from its starting temperature. How many degrees did it drop ?

OpenStudy (anonymous):

wait can you rephrase that in another way?

OpenStudy (phi):

what was the temperature of the metal before we put it in the water ?

OpenStudy (anonymous):

the temp was 100?

OpenStudy (phi):

yes, the metal starts at 100C and drops to 31.7 C ΔT= 100-31.7

OpenStudy (anonymous):

I feel like crying right now, you had to fight and claw your way into my brain so that I can understand thank you!

OpenStudy (anonymous):

ok so now I have to subtract 100-31.7 to get the C?

OpenStudy (phi):

almost. ΔT= 100-31.7 remember we are doing 728.156= 27.776 x c x ΔT

OpenStudy (anonymous):

ok so the T= 68.3

OpenStudy (anonymous):

728.156= 27.776 x c x 68.3

OpenStudy (phi):

and the equation (so far) is 728.156= 27.776 x c x 68.3 we can simplify the right side by multiplying 27.776 x 68.3

OpenStudy (phi):

you should get 728.156= 1897.1 c (we can leave out the x, we know we are multiplying) to find c, we divide both sides by 1897.1 like this: 728.156/1897.1 = c * 1897.1/1897.1 which simplifies to 728.156/1897.1 = c or c = 728.156/1897.1

OpenStudy (anonymous):

I got it, but don't go too fast on me like that or I could get lost. Um so C= 0.38?

OpenStudy (phi):

yes, c = 0.38 J / (g × °C) that means it might be copper or zinc

OpenStudy (anonymous):

which part means it might be copper or zinc, the (g x C)?

OpenStudy (phi):

the units are Joules (that is energy) per gram (how much metal) per degree we can try to match 0.38 J / (g × °C) to a table of specific heat capacity of different metals, and the closest I can find are copper and zinc which both have 0.39 J / (g × °C)

OpenStudy (anonymous):

oh ok that would make sense since this is a zinc table

OpenStudy (phi):

now for you unknown metal

OpenStudy (anonymous):

phew ok I think I understand now for the most part, let me show you I get it . Here's the table, I'll show you what to do. 25.605 g-Measured mass of metal 24 mL- Distilled water measurement 25 C- Distilled water temp 100 C- Temp of metal 0 C- temp of mixture

OpenStudy (anonymous):

Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qunknown metal Using the formula qunknown metal = m × c × ΔT, calculate the specific heat of the metal.

OpenStudy (phi):

I have a problem with your data. If we start with water at 25C and add metal at 100C the mixture CAN NOT go to 0º C. The water should get hotter than 25 and the metal should cool down (but not to 0)

OpenStudy (anonymous):

when I did the lab that's what I got, do you want to re do the lab on that part?

OpenStudy (phi):

I think we better.

OpenStudy (anonymous):

ok hold on I'm doing that right now give me 2 min

OpenStudy (anonymous):

@phi : i got 28.7 C this time around

OpenStudy (phi):

OK, can you do the water part? qwater = m × c × ΔT

OpenStudy (anonymous):

um, first we find the joules right?

OpenStudy (phi):

yes, but the first thing is write down m (mass in grams) c (4.18) and ΔT (28.7 C - starting temp of water)

OpenStudy (anonymous):

quick question where is the 4.18 coming from again?

OpenStudy (phi):

it is the known c for water.

OpenStudy (anonymous):

25.605 x 4.18 x (100-28.7)

OpenStudy (phi):

no. first we do the equation for the WATER. how much water do you have? what temperature did it start at ?

OpenStudy (anonymous):

24 mL?

OpenStudy (phi):

I did not do the experiment. When you just did it, how much water did you use ?

OpenStudy (anonymous):

hold on, this is a virtual lab

OpenStudy (phi):

yes, I assume this is not real or it would take a lot longer to do.

OpenStudy (anonymous):

yea, btw thanks for having a lot of patience with me, this is not my thing esp when it involves math. Reading and English is my thing so when you need help it could be like a collab

OpenStudy (anonymous):

but hold on

OpenStudy (anonymous):

i used 25 g of water when I first start.

OpenStudy (phi):

can you list all the data right now?

OpenStudy (phi):

ok. now we use qwater = m × c × ΔT m stands for the mass of the water. They give the water in ml, but water is 1 g per ml so that means we can just use the number they give us. c is 4.18 (because we are told that) ΔT is the difference between the temp of the water and the temp of the mixture. can you fill in the numbers?

OpenStudy (anonymous):

what do you mean for the number they give us?

OpenStudy (phi):

how much water is used in the experiment ?

OpenStudy (anonymous):

rephrase that

OpenStudy (anonymous):

cause I'm confused when you say that

OpenStudy (phi):

how much distilled water is used in the experiment?

OpenStudy (anonymous):

measurement or temperature?